# Instability and breaking of internal waves in a horizontal shear layer

### v1

@@ 2.2.1 Trapping and turning levels
Given $U(y)=\Delta u\tanh(y/h)$, it follows from (15) that $k_{y}$ increases or decreases monotonically for $k_{x}<0$ or $k_{x}>0$, respectively. We assume henceforth that $k_{y}>0$ initially, so that $k_{x}<0$ means $k_{y}>0$ for all times. In particular, the situation may arise where $k_{y}\to\infty$, in which case the intrinsic frequency approaches the buoyancy frequency $\Omega\to 1$ (corresponding to $\Omega^{*}\to N^{*}$ in dimensional variables). The location $y=y_{t}$ where $\Omega=1$ (when it exists) is referred to as a trapping level, where the group velocity in the $y$-direction approaches zero asymptotically. (Formally however, the WKBJ theory becomes invalid before the trapping level is reached.) It follows from the Hamiltonian structure of (14) and (15) that $\omega$ is constant along rays (e.g. sutherland2010internal; buhler2014waves). Then, equating the initial value of $\omega=\omega_{0}$ (for which the intrinsic frequency $\Omega=\Omega_{0})$ with the value at the trapping level determines $y_{t}$ implicitly as
@@
where $y_{0}$ is the initial location of the wave packet. A detailed description of the asymptotic approach to $y_{t}$ is provided by badulin1985trapping. The trapping level exists only when the right hand side of equation (16) falls between 1-1 and $1$. Assuming without loss of generality that the wave packet propagates in the positive $y$ direction and is also located initially outside of the shear layer such that $\tanh(y_{0}/h)\approx-1$, it follows that we must have $0\leq(\Omega_{0}-1)/(\Delta uk_{x})\leq 2$ for the trapping level to exist.
When $k_{x}>0$, $k_{y}$ decreases and internal reflection may occur where $k_{y}$ passes through 0 and becomes negative. However, the WKBJ analysis fails for small $k_{y}$ and a separate solution must be obtained in the vicinity of the turning level or an alternative approach used (e.g. bakas2009gravitya). The location of the turning level where $k_{y}=0$ (when it exists) may be obtained from an equation analogous to (16).
@@ 2.2.2 Kinetic and potential energy partition
It follows from equations (13) and (17) that wave advection of momentum can result in ‘excess’ kinetic energy contained purely in the streamwise velocity component, here referring to energy not accounted for by the usual wave energy density $s^{2}/2k_{z}^{2}$. The presence of this additional energy has no effect on the dynamics of wave propagation (for example, the group velocity and dispersion relation) at leading order because $Uk_{x}\sim O(1)$. In particular, this applies to the conservation equation for wave action $\mathcal{A}$ (e.g. bretherton1968wavetrains), which for the case of a wave packet periodic in $x$ and $z$ reads
@@
where $\mathcal{A}\equiv s^{2}/(2k_{z}^{2})/\Omega$. Here we have also included a viscous term on the right hand side due to grimshaw1974internal (see also koop1981preliminary; howland2021shear).
The discussion above suggests that perturbation energy may change due to two distinct mechanisms. Wave advection of momentum leads to a change in the kinetic energy contained entirely in the streamwise velocity component according to the local polarisation relations (13). On the other hand, wave refraction by the horizontal shear modifies the intrinsic frequency and group velocity and hence changes the local wave steepness $s$ via wave action conservation. In the context of the ray-tracing equations, these behaviours may be interpreted as local manifestations of the global transient energy growth mechanisms referred to by bakas2009gravitya; bakas2009gravityb as growth due to the roll (lift-up) mechanism and to downgradient Reynolds stresses, respectively. Note that these mechanisms operate independently of one another when $Uk_{x}\sim O(1)$, but may be simultaneously active.
@@ 2.2.3 Limiting dynamics and instability
As noted above, when $k_{x}=0$ it is clear from (15) that the local wave state is unaffected by the shear flow. Thus in this case energy growth occurs purely through wave advection of mean flow momentum. As long as perturbations remain independent of $x$, the total streamwise momentum $U(y)+u(y,z,t)$ behaves simply as a passive scalar in the system. This behaviour is often referred to as the ‘lift-up’ mechanism for shear flow perturbations dating back to the works of landahl1980note; ellingsen1975stability; moffat1967interaction and leads to the generation of $y$-vorticity and hence vertical shear via vortex tilting (butler1992three). Note that $\Omega^{2}=k_{y}^{2}/|\mathbf{k}|^{2}$ in this case so that
@@
Hence the ratio of perturbation energy to wave energy is $1+Fr_{h}^{2}k_{z}^{2}/k_{y}^{2}$, where $Fr_{h}=\partial U/\partial y=(\partial U^{*}/\partial y^{*})/N^{*}$ is the local horizontal Froude number associated with the background shear flow. Though the analysis here is strictly local, it is worth pointing out that this is an equivalent scaling to that derived for the optimal maximum total energy growth for $x$-independent perturbations in an unbounded uniform horizontal shear flow by bakas2009gravitya. The greatest perturbation energies result from waves with large horizontal to vertical aspect ratio $k_{z}/k_{y}$ in a flow with large $Fr_{h}$.
@@
describing the instantaneous proportion of excess kinetic energy due to wave advection of momentum. Note that $F$ varies along rays due to changes in $k_{y}$ and $\partial U/\partial y$. The evolution may be readily calculated by solving the ray tracing equations (15) numerically. badulin1985trapping demonstrate that the wave steepness $s$ increases due to energy extraction from the shear near trapping levels. As suggested by bakas2009gravitya, when $k_{x}<0$ (so that $k_{y}$ is increasing) and $F$ is small so that wave advection of momentum is relatively unimportant, corresponding amplitude growth may lead to the formation of regions of unstable buoyancy gradients and hence the potential for convective instability. As $F$ increases, the importance of wave advection of momentum increases, which leads to a potentially significant increase in the vertical shear. This can be seen in our analysis by using the local wave polarisation relations (13) to compute that $|\partial u/\partial z|^{2}+|\partial v/\partial z|^{2}=k_{z}^{2}/|\mathbf{k}|^{2}+F^{2}$.
@@
defined here based on a locally parallel shear flow $u_{\beta}=u\sin\beta+v\cos\beta$ obtained from the projection of the horizontal velocity field onto a chosen azimuthal direction $\beta$ in which unstable perturbations are assumed to grow. A criterion of $Ri_{g}<1/4$ somewhere in the flow is often adopted as an indicator of the potential for shear instability based on the classical Miles-Howard theorem (miles1961stability; howard1961note), though the formal assumptions required for the theorem to be applicable are invalid even for the simple linear wave state considered here. Nonetheless, $Ri_{g}<1/4$ has proven useful as a qualitative indicator of flow instability in a wide range of flows. For a linear internal gravity wave given by the polarisation relations (5), it may be shown (e.g. thorpe1999breaking) that, for statically stable configurations $s<1$, the minimum value of $Ri_{g}$ is $1/2$, and so shear instability is not expected. However, when the wave interacts with a horizontal shear flow such that the streamwise velocity field is modified according to equations (13), significant changes to $Ri_{g}$ may be expected.
In the limit $k_{x}=0$, the horizontal velocity components $u$ and $v$ differ by a phase of $\pi/2$. In this case, an analogy with inertia-gravity waves in rotating stratified flows provides an interesting point of comparison. In particular, the polarisation between these two components is the same as a linear inertia-gravity wave propagating in the $y$-$z$ plane under the association $\partial U/\partial y\leftrightarrow-f$ (importantly however, the same is not true of the dispersion relation). Unlike those in systems with $f=0$, such waves may exhibit $Ri_{g}<1/4$ for $s<1$, with the resulting dynamics closely resembling classical parallel shear instability (lelong1998inertiaa). The differences between the resulting energetics and those associated with convective instabilities may be significant (thorpe2018models). For $F\neq 0$, the local gradient Richardson number for the horizontally sheared system considered here is calculated to be
@@
where $\phi=\mathbf{k}\cdot\mathbf{x}-\Omega t$ is the local wave phase and $R=(\partial U/\partial y)/\Omega$. When $k_{x}=0$, as for the plane inertia gravity wave studied in lelong1998inertiaa, $Ri_{g}$ attains its absolute minimum value at $\beta=\pi/2$ and $\phi=\pi$, corresponding to perturbations transverse to the direction of wave propagation. This minimum is $(1-s)/F^{2}s^{2}$, suggesting that $F$ may be useful as a proxy for shear instability. Indeed, because $k_{z}/k_{y}$ represents a horizontal to vertical wave aspect ratio, the expression (20) takes the form of an inverse Richardson number. For non-zero $k_{x}$, the values of $\beta$ and $\phi$ minimising (22) remain close to $\pi/2$ and $\pi$ as long as $(k_{x}/k_{y})/R\lesssim 1$, which does not turn out to be an important restriction for the cases considered here.
The physical meaning of $F$ breaks down when $k_{y}$ becomes small such that the WKBJ approximation is no longer valid. However, for rapid growth in the streamwise velocity field (i.e. large $\partial U/\partial y$), the small-amplitude linear approximation may break down before this point. In the special case $k_{x}=0$, Lagrangian conservation of total streamwise momentum limits the maximum amplitude of the streamwise velocity to be at most $\Delta u$ and the expression for $F$ is certainly meaningless beyond this limit. With these restrictions in mind, we consider the ray tracing equations under some representative cases to illustrate the variability in dynamics that can occur due to the influence of horizontal shear. We fix initial values of $k_{y}=0.25$ and $k_{z}=1$ (consistent with the direct numerical simulations described in section 3) and select values of $k_{x}$ and $\partial U/\partial y>0$ to produce scenarios where the wave is trapped ($k_{x}<0$), reflected ($k_{x}>0$) and unaffected by the shear ($k_{x}=0$). The ray tracing equations (14) and (15) are solved numerically alongside equation (18) governing the evolution of wave action $\mathcal{A}$. Local values of the wave steepness $s$ are computed from $s^{2}=2\mathcal{A}\Omega k_{z}^{2}$.
@@
We first consider the behaviour of $s$. When $k_{x}<0$, $k_{y}$ increases monotonically along rays, leading to a decrease in the group velocity as the trapping level at $y\approx 0$ is approached. This leads to the significant increase in the wave steepness observed in panels $(a)$ and $(b)$ before viscosity acts to damp further energy growth, consistent with the results of badulin1985trapping. When $k_{x}=0$, no refraction takes place and wave energy is conserved along rays in the inviscid limit. This corresponds to the approximately constant value of $s$ seen in panel $(c)$, which decreases only due to viscosity. When $k_{x}=0.05$ in panel $(d)$ such that a turning level exists at $y\approx 0$, $s$ similarly remains approximately constant as the turning level is approached, until the linear WKBJ approximation becomes invalid at $y\approx-2$. The parameter $F$ exhibits distinctly different behaviour to that of $s$. When $k_{x}<0$, the growth in $F$ depends on $\partial U/\partial y$. When $\Delta u/h=0.125$ is small, as in panel $(a)$, the dynamics are dominated by wave refraction and $F$ remains close to 0. For larger $\Delta u/h=1$ (panel $b$), $F$ is able to grow to $O(1)$ before growth in $k_{y}$ eventually dominates. When $k_{x}=0$ (panel $c$), energy growth is due entirely to the lift-up mechanism, demonstrated by the increase in $F$ as the wave approaches the center of the shear layer. A similar behaviour is observed for $k_{x}>0$ (panel $d$). Note that, for a given $k_{x0}$, $k_{z0}$ and $\Delta u/h$, the maximum value of $F$ is obtained when $k_{y}^{2}=k_{x0}^{2}<k_{y0}^{2}$, explaining why a larger maximum is predicted for the reflected case where $k_{y}$ may vary.
@@ 3 Numerical simulations
The linear theory in the previous section suggests that a range of wave-breaking behaviours are possible in a horizontal shear layer, depending on the initial flow parameters. However, this analysis is at best heuristic, being strictly valid only for small-amplitude wave perturbations in the WKBJ approximation. To further understand the flow evolution, we perform a series of complementary fully nonlinear direct numerical simulations (DNS) intended to span the range of predicted breaking behaviours. Simulations are performed using the multi-parallel pseudo-spectral channel flow solver Diablo (Taylor08), which employs a second-order finite difference discretization in the wall-bounded direction, here taken to be in the direction of horizontal shear. The code uses a third-order mixed implicit/explicit Runge-Kutta/Crank-Nicolson scheme for time marching, with a 2/3 dealiasing rule applied to the nonlinear terms.
@@ 3.1 Overview
Under the expectation that $k_{z}$ remains constant throughout the wave-shear interaction, it is convenient to take $K^{*}=k^{*}_{z0}$ in the non-dimensionalisation outlined in §2.1. In this way, the dimensionless initial wavenumber vector $\mathbf{k}_{0}=(k_{x0},k_{y0},1)$ has corresponding frequency $\Omega_{0}$ satisfying $\Omega_{0}^{2}=(k_{x0}^{2}+k_{y0}^{2})/(1+k_{x0}^{2}+k_{y0}^{2})$, and the initial (maximum) wave energy density $E_{0}=s_{0}^{2}/(2k_{z0}^{2})=s_{0}^{2}/2$. Most of the simulations presented here have $k_{y0}=0.25$, with values of $k_{x0}$ then selected to span a range of predicted trapping and reflecting scenarios. The dimensionless initial frequency $0.125<\Omega_{0}<0.38$ (corresponding to propagation angles of $13^{\circ}$-$21^{\circ}$ between $\mathbf{k}_{0}$ and the horizontal) is large enough such that the neglect of planetary rotation is reasonable (under the assumption $f\ll N$) without being so large that waves quickly become unstable to a parametric instability (lombard1996breakdown). We note these values are within the estimated range of high frequency waves observed in the ocean thermocline by alford2000observations. The packet width scale is chosen as $a=6\pi$, with the envelope centered at $y_{0}=-2a$.

### v2

@@ 2.2.1 Trapping and turning levels
Given $U(y)=\Delta u\tanh(y/h)$, it follows from (15) that $k_{y}$ increases or decreases monotonically for $k_{x}<0$ or $k_{x}>0$, respectively. We assume henceforth that $k_{y}>0$ initially, so that $k_{x}<0$ means $k_{y}>0$ for all times. In particular, the situation may arise where $k_{y}\to\infty$, in which case the intrinsic frequency approaches the buoyancy frequency $\Omega\to 1$ (corresponding to $\Omega^{*}\to N^{*}$ in dimensional variables). The location $y=y_{t}$ where $\Omega=1$ (when it exists) is referred to as a trapping level, where the group velocity in the $y$-direction approaches zero asymptotically. (Formally however, the WKBJ theory becomes invalid before the trapping level is reached.) It follows from the Hamiltonian structure of (14) and (15) that $\omega$ is constant along rays (e.g. Sutherland, 2010; Bühler, 2014). Then, equating the initial value of $\omega=\omega_{0}$ (for which the intrinsic frequency $\Omega=\Omega_{0})$ with the value at the trapping level determines $y_{t}$ implicitly as
@@
where $y_{0}$ is the initial location of the wave packet. A detailed description of the asymptotic approach to $y_{t}$ is provided by Badulin et al. (1985). The trapping level exists only when the right hand side of equation (16) falls between 1-1 and $1$. Assuming without loss of generality that the wave packet propagates in the positive $y$ direction and is also located initially outside of the shear layer such that $\tanh(y_{0}/h)\approx-1$, it follows that we must have $0\leq(\Omega_{0}-1)/(\Delta uk_{x})\leq 2$ for the trapping level to exist.
When $k_{x}>0$, $k_{y}$ decreases and internal reflection may occur where $k_{y}$ passes through 0 and becomes negative. However, the WKBJ analysis fails for small $k_{y}$ and a separate solution must be obtained in the vicinity of the turning level or an alternative approach used (e.g. Bakas and Farrell, 2009a). The location of the turning level where $k_{y}=0$ (when it exists) may be obtained from an equation analogous to (16).
@@ 2.2.2 Kinetic and potential energy partition
It follows from equations (13) and (17) that wave advection of momentum can result in ‘excess’ kinetic energy contained purely in the streamwise velocity component, here referring to energy not accounted for by the usual wave energy density $s^{2}/2k_{z}^{2}$. The presence of this additional energy has no effect on the dynamics of wave propagation (for example, the group velocity and dispersion relation) at leading order because $Uk_{x}\sim O(1)$. In particular, this applies to the conservation equation for wave action $\mathcal{A}$ (e.g. Bretherton and Garrett, 1968), which for the case of a wave packet periodic in $x$ and $z$ reads
@@
where $\mathcal{A}\equiv s^{2}/(2k_{z}^{2})/\Omega$. Here we have also included a viscous term on the right hand side due to Grimshaw (1974) (see also Koop, 1981; Howland et al., 2021).
The discussion above suggests that perturbation energy may change due to two distinct mechanisms. Wave advection of momentum leads to a change in the kinetic energy contained entirely in the streamwise velocity component according to the local polarisation relations (13). On the other hand, wave refraction by the horizontal shear modifies the intrinsic frequency and group velocity and hence changes the local wave steepness $s$ via wave action conservation. In the context of the ray-tracing equations, these behaviours may be interpreted as local manifestations of the global transient energy growth mechanisms referred to by Bakas and Farrell (2009a, b) as growth due to the roll (lift-up) mechanism and to downgradient Reynolds stresses, respectively. Note that these mechanisms operate independently of one another when $Uk_{x}\sim O(1)$, but may be simultaneously active.
@@ 2.2.3 Limiting dynamics and instability
As noted above, when $k_{x}=0$ it is clear from (15) that the local wave state is unaffected by the shear flow. Thus in this case energy growth occurs purely through wave advection of mean flow momentum. As long as perturbations remain independent of $x$, the total streamwise momentum $U(y)+u(y,z,t)$ behaves simply as a passive scalar in the system. This behaviour is often referred to as the ‘lift-up’ mechanism for shear flow perturbations dating back to the works of Landahl (1980); Ellingsen and Palm (1975); Moffat (1967) and leads to the generation of $y$-vorticity and hence vertical shear via vortex tilting (Butler and Farrell, 1992). Note that $\Omega^{2}=k_{y}^{2}/|\mathbf{k}|^{2}$ in this case so that
@@
Hence the ratio of the bracketed terms in equation (19) is $1+Fr_{h}^{2}k_{z}^{2}/k_{y}^{2}$, where $Fr_{h}=\partial U/\partial y=(\partial U^{*}/\partial y^{*})/N^{*}$ is the local horizontal Froude number associated with the background shear flow. Though the analysis here is strictly local, it is worth pointing out that this is an equivalent scaling to that derived for the optimal maximum total energy growth for $x$-independent perturbations in an unbounded uniform horizontal shear flow by Bakas and Farrell (2009a). The greatest perturbation energies result from waves with large horizontal to vertical aspect ratio $k_{z}/k_{y}$ in a flow with large $Fr_{h}$.
@@
describing the instantaneous proportion of excess kinetic energy due to wave advection of momentum. Note that $F$ varies along rays due to changes in $k_{y}$ and $\partial U/\partial y$. The evolution may be readily calculated by solving the ray tracing equations (15) numerically. Badulin et al. (1985) demonstrate that the wave steepness $s$ increases due to energy extraction from the shear near trapping levels. As suggested by Bakas and Farrell (2009a), when $k_{x}<0$ (so that $k_{y}$ is increasing) and $F$ is small so that wave advection of momentum is relatively unimportant, corresponding amplitude growth may lead to the formation of regions of unstable buoyancy gradients and hence the potential for convective instability. As $F$ increases, the importance of wave advection of momentum increases, which leads to a potentially significant increase in the vertical shear. This can be seen in our analysis by using the local wave polarisation relations (13) to compute that $|\partial u/\partial z|^{2}+|\partial v/\partial z|^{2}=k_{z}^{2}/|\mathbf{k}|^{2}+F^{2}$.
@@
defined here based on a locally parallel shear flow $u_{\beta}=u\sin\beta+v\cos\beta$ obtained from the projection of the horizontal velocity field onto a chosen azimuthal direction $\beta$ in which unstable perturbations are assumed to grow. A criterion of $Ri_{g}<1/4$ somewhere in the flow is often adopted as an indicator of the potential for shear instability based on the classical Miles-Howard theorem (Miles, 1961; Howard, 1961), though the formal assumptions required for the theorem to be applicable are invalid even for the simple linear wave state considered here. Nonetheless, $Ri_{g}<1/4$ has proven useful as a qualitative indicator of flow instability in a wide range of flows. For a linear internal gravity wave given by the polarisation relations (5), it may be shown (e.g. Thorpe, 1999) that, for statically stable configurations $s<1$, the minimum value of $Ri_{g}$ is $1/2$, and so shear instability is not expected. However, when the wave interacts with a horizontal shear flow such that the streamwise velocity field is modified according to equations (13), significant changes to $Ri_{g}$ may be expected.
In the limit $k_{x}=0$, the horizontal velocity components $u$ and $v$ differ by a phase of $\pi/2$. In this case, an analogy with inertia-gravity waves in rotating stratified flows provides an interesting point of comparison. In particular, the polarisation between these two components is the same as a linear inertia-gravity wave propagating in the $y$-$z$ plane under the association $\partial U/\partial y\leftrightarrow-f$ (importantly however, the same is not true of the dispersion relation). Unlike those in systems with $f=0$, such waves may exhibit $Ri_{g}<1/4$ for $s<1$, with the resulting dynamics closely resembling classical parallel shear instability (Lelong and Dunkerton, 1998a). The differences between the resulting energetics and those associated with convective instabilities may be significant (Thorpe, 2018). For $F\neq 0$, the local gradient Richardson number for the horizontally sheared system considered here is calculated to be
@@
where $\phi=\mathbf{k}\cdot\mathbf{x}-\Omega t$ is the local wave phase and $R=(\partial U/\partial y)/\Omega$. It is demonstrated in Appendix A that the minimum value of $Ri_{g}$ is very close to (and in some cases exactly equal to) $(1-s)/F^{2}s^{2}$ when this quantity is smaller than $0.5$. This value is achieved at $\beta=\pi/2$ and $\phi=\pi$, corresponding to perturbations transverse to the direction of wave propagation (c.f. Lelong and Dunkerton, 1998a). It is therefore apparent that $F$ may be useful as a proxy for shear instability. Indeed, because $k_{z}/k_{y}$ represents a horizontal to vertical wave aspect ratio, the expression (20) takes the form of an inverse Richardson number.
The physical meaning of $F$ breaks down when $k_{y}$ becomes small such that the WKBJ approximation is no longer valid. However, for rapid growth in the streamwise velocity field (i.e. large $\partial U/\partial y$), the small-amplitude linear approximation may break down before this point. In the special case $k_{x}=0$, Lagrangian conservation of total streamwise momentum limits the maximum amplitude of the streamwise velocity to be at most $\Delta u$ and the expression for $F$ is certainly meaningless beyond this limit. With these restrictions in mind, we consider the ray tracing equations under some representative cases to illustrate the variability in dynamics that can occur due to the influence of horizontal shear. We fix $k_{z}=1$ and choose values of $k_{x}$, $k_{y}$, $\Delta u$ and $h$ to produce scenarios where the wave is trapped ($k_{x}<0$), reflected ($k_{x}>0$) and unaffected by the shear ($k_{x}=0$), consistent with the direct numerical simulations described in section 3. The ray tracing equations (14) and (15) are solved numerically alongside equation (18) governing the evolution of wave action $\mathcal{A}$. Local values of the wave steepness $s$ are computed from $s^{2}=2\mathcal{A}\Omega k_{z}^{2}$.
@@
We first consider the behaviour of $s$. When $k_{x}<0$, $k_{y}$ increases monotonically along rays, leading to a decrease in the group velocity as the trapping level at $y\approx 0$ is approached. This leads to the significant increase in the wave steepness observed in panels $(a)$ and $(b)$ before viscosity acts to damp further energy growth, consistent with the results of Badulin et al. (1985). When $k_{x}=0$, no refraction takes place and wave energy is conserved along rays in the inviscid limit. This corresponds to the approximately constant value of $s$ seen in panel $(c)$, which decreases only due to viscosity. When $k_{x}=0.05$ in panel $(d)$ such that a turning level exists at $y\approx 0$, $s$ similarly remains approximately constant as the turning level is approached, until the linear WKBJ approximation becomes invalid at $y\approx-2$. The parameter $F$ exhibits distinctly different behaviour to that of $s$. When $k_{x}<0$, the growth in $F$ depends on $\partial U/\partial y$. When $\Delta u/h=0.125$ is small, as in panel $(a)$, the dynamics are dominated by wave refraction and $F$ remains close to 0. For larger $\Delta u/h=1$ (panel $b$), $F$ is able to grow to $O(1)$ before growth in $k_{y}$ eventually dominates. When $k_{x}=0$ (panel $c$), energy growth is due entirely to the lift-up mechanism, demonstrated by the increase in $F$ as the wave approaches the center of the shear layer. A similar behaviour is observed for $k_{x}>0$ (panel $d$). Note that, for a given $k_{x0}$, $k_{z0}$ and $\Delta u/h$, the maximum value of $F$ is obtained when $k_{y}^{2}=k_{x0}^{2}<k_{y0}^{2}$, explaining why a larger maximum is predicted for the reflected case where $k_{y}$ may vary.
@@ 3 Numerical simulations
The linear theory in the previous section suggests that a range of wave-breaking behaviours are possible in a horizontal shear layer, depending on the initial flow parameters. However, this analysis is at best heuristic, being strictly valid only for small-amplitude wave perturbations in the WKBJ approximation. To further understand the flow evolution, we perform a series of complementary fully nonlinear direct numerical simulations (DNS) intended to span the range of predicted breaking behaviours. Simulations are performed using the multi-parallel pseudo-spectral channel flow solver Diablo (Taylor, 2008), which employs a second-order finite difference discretization in the wall-bounded direction, here taken to be in the direction of horizontal shear. The code uses a third-order mixed implicit/explicit Runge-Kutta/Crank-Nicolson scheme for time marching, with a 2/3 dealiasing rule applied to the nonlinear terms.
@@ 3.1 Overview
Under the expectation that $k_{z}$ remains constant throughout the wave-shear interaction, it is convenient to take $K^{*}=k^{*}_{z0}$ in the non-dimensionalisation outlined in §2.1. In this way, the dimensionless initial wavenumber vector $\mathbf{k}_{0}=(k_{x0},k_{y0},1)$ has corresponding frequency $\Omega_{0}$ satisfying $\Omega_{0}^{2}=(k_{x0}^{2}+k_{y0}^{2})/(1+k_{x0}^{2}+k_{y0}^{2})$, and the initial (maximum) wave energy density $E_{0}=s_{0}^{2}/(2k_{z0}^{2})=s_{0}^{2}/2$. Most of the simulations presented here have $k_{y0}=0.25$, with values of $k_{x0}$ then selected to span a range of predicted trapping and reflecting scenarios. The dimensionless initial frequency $0.125<\Omega_{0}<0.38$ (corresponding to propagation angles of $13^{\circ}$-$21^{\circ}$ between $\mathbf{k}_{0}$ and the horizontal) is large enough such that the neglect of planetary rotation is reasonable (under the assumption $f\ll N$) without being so large that waves quickly become unstable to a parametric instability (Lombard and Riley, 1996b). We note these values are within the estimated range of high frequency waves observed in the ocean thermocline by Alford and Pinkel (2000). The packet width scale is chosen as $a=6\pi$, with the envelope centered at $y_{0}=-2a$. Simulations are initialised with $\mathbf{u}(\mathbf{x},t=0)=\mathbf{U}(y)+\mathbf{u}_{\mathrm{wave}}(\mathbf{x})$ and $b=b_{\mathrm{wave}}(\mathbf{x})$ from (7) and (8). Additionally, white noise of amplitude $10^{-5}$ is added to each component of the velocity field in order to trigger instabilities leading to turbulence.